# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v3

@@ 5 Mitigating Reward Hacking with Occupancy Measure Regularization
We now discuss how our new definition of reward hacking motivates better methods for preventing it. Ideally, we would like to be able to optimize a proxy reward function and have it translate into an improvement in true reward over the base policy. To understand how this might be possible, consider again the traffic environment. The reward hacking policy exhibits behavior which is very unlikely under the base policy: human drivers hardly ever stop indefinitely on on-ramps. That is, optimizing the proxy reward leads to out-of-distribution states and actions where the correlation that made the proxy good in the first place breaks down. Visually, this can be seen in the second row of Figure 2: the reward hacking policy usually finds state-action pairs with high proxy reward and low true reward that were very unlikely to be reached by the base policy.
Thus, one solution to prevent reward hacking could be to optimize the proxy reward while avoiding states that are unlikely under the base policy. The following theorem formalizes this idea.
@@
Suppose that $\tilde{R}$ is an $r$-correlated-proxy for the true reward function $R$, and let $\sigma_{\tilde{R}}$ and $\sigma_{R}$ be defined as in Definition 4.1. Then for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$ (i.e., $\mu_{{\pi_{\text{base}}}}(s,a)=0\Rightarrow\mu_{\pi}(s,a)=0$), we have
$\frac{J(\pi,R)-J({\pi_{\text{base}}},R)}{\sigma_{R}}\geq\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right),$ (4)
where $\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)=\mathbb{E}_{\mu_{\pi}}\left[\frac{\mu_{\pi}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}-1\right]$ is the $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{base}}}$.
See Appendix A for the proof. Equation (10) gives a lower bound on how much the policy $\pi$ improves over the base policy $\pi_{\text{base}}$} in terms of the true reward, normalized by the standard deviation of the true reward under the base policy. Although this is exactly what we would like to maximize, we must instead optimize the right-hand side (RHS) of (10) since we cannot access the true reward directly. The RHS consists of two terms: one measuring the normalized improvement of the policy $\pi$’s proxy reward over the base policy and the other penalizing the divergence of $\pi$’s occupancy measure from that of $\pi_{\text{base}}$}. Optimizing the first term alone often leads to reward hacking, but by both incentivizing $\pi$ to achieve high proxy reward and stay close to the base policy, $\pi$ can achieve high true reward.
@@
$\text{maximize}\quad J(\pi,\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\qquad\text{where}\qquad\lambda=\sigma_{\tilde{R}}\sqrt{1-r^{2}}.$ (5)
The amount of regularization needed to improve on the base policy depends on the strength of correlation $r$. The higher the correlation, the lower the regularization strength $\sqrt{1-r^{2}}$. Given the prefactor of $\frac{1}{r}$ on the RHS of (10), it may appear that a lower correlation leads to a larger gain in true reward. However, Lemma A.2 shows that in fact, the lower bound decreases as a function of $r$.
@@
$\text{maximize}\quad J(\pi,\tilde{R})-\lambda(1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{\text{KL}}\big(\pi(\cdot\mid s_{t})\;\|\;\mu_{\pi_{\text{base}}}(\cdot\mid s_{t})\big)\right].$ (6)
That is, in RLHF, the expected KL divergence between the action distributions of $\pi$ and $\pi_{\text{base}}$} is penalized. Action distribution divergence is easy to calculate and optimize, and training LLMs with (6) seems to work well in practice. However, unlike our objective in (5), (6) lacks theoretical guarantees, and it is unclear if it works in other environments.
@@
Figure 4: Our theory also suggests reward hacking can more effectively be prevented by regularizing with $\chi^{2}$ divergence instead of KL divergence. This plot illustrates how $\chi^{2}$ regularization is more effective at preventing reward hacking in RLHF. Both divergences can be written as the expectation of a function $g\big(\log\big(\mu_{\pi}(s,a)/\mu_{{\pi_{\text{base}}}}(s,a)\big)\big)$ which increases the penalty on state-action pairs based on how far the log-ratio is from zero. The $g(\cdot)$ associated with KL divergence only increases slowly for large log-ratios, so policies trained with KL divergence may produce nonsensical text. In contrast, the $g(\cdot)$ for $\chi^{2}$ divergence increases exponentially, better constraining the LLM to produce text similar to the SFT policy.
Occupancy measure vs. action distribution regularization.  While AD regularization works well for RLHF, this may be because RLHF is essentially a contextual bandit problem, meaning that $\gamma=0$; in this case, OM and AD divergence are equivalent (see Appendix A.3). However, in other cases, AD divergence may not suffice to prevent reward hacking behavior. This is because, in longer-horizon environments, a small change in action distribution at a single state can lead to a much higher probability of reaching undesirable states. Figure 3 shows an example of this in the glucose monitoring environment: policies that are close in action distribution regularization produce vastly different patient glucose levels. Occupancy measure regularization avoids this issue by directly preventing the distribution of glucose levels from differing too much from the base policy. In Theorem A.5 in the appendix, we show that in general it is impossible to lower bound the improvement in true reward using almost any form of AD regularization, in contrast to our results on OM regularization.
@@
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{{\pi_{\text{base}}}})=\mathbb{E}_{\mu_{\pi}}\left[d(s,a)+e^{-d(s,a)}\right]\qquad\chi^{2}(\mu_{\pi}\;\|\;\mu_{{\pi_{\text{base}}}})=\mathbb{E}_{\mu_{\pi}}\left[e^{d(s,a)}+e^{-d(s,a)}\right]$
$\text{where}\qquad d(s,a)=\log\big(\mu_{\pi}(s,a)/\mu_{{\pi_{\text{base}}}}(s,a)\big).$ (7)
As $d(s,a)$ increases, the optimized policy is visiting state-action pairs that are less likely under the base policy. However, KL divergence only penalizes increases in $d(s,a)$ linearly, while $\chi^{2}$ penalizes them exponentially, resulting in stronger regularization even with a lower coefficient. Figure 4 plots the functions in (7) and shows how in practice $\chi^{2}$ divergence better prevents reward hacking in RLHF.
@@
Table 1: We compare using various types of regularization to prevent reward hacking in the five environments form Figure 2. The median true reward and standard deviation across 5 random seeds is shown for the best regularization coefficient for each type of regularization. The bottom rows show results for the baselines: the base policy $\pi_{\text{base}}$}, a policy trained on the proxy reward without regularization (exhibiting reward hacking), and a policy trained on the true reward function (impossible in practice, but included as an upper bound on performance). We find that occupancy measure regularization consistently improves on action distribution regularization, and that $\chi^{2}$ divergence often outperforms KL divergence.
@@ 6 Experiments
$\text{and}\qquad\phi$ $=\arg\min_{\phi^{\prime}}\;\mathbb{E}_{\mu_{\pi}}\Big[\,\log(1+e^{-\hat{d}_{\phi^{\prime}}(s,a)})\,\Big]+\,\mathbb{E}_{\mu_{\pi_{\text{base}}}}\Big[\,\log(1+e^{\hat{d}_{\phi^{\prime}}(s,a)})\,\Big].$ (8)
That is, optimizing a discriminator network $\hat{d}_{\phi}$ to minimize the given loss can be used to estimate and optimize $\chi^{2}$ OM divergence. We alternately train $\hat{d}_{\phi}$ via gradient descent on (8) and the policy $\pi$ via PPO based on the adjusted reward. We call this algorithm Occupancy-Regularized Policy Optimization (ORPO). See Appendix B for a full derivation of the approximations used in ORPO and Algorithm 1 for a formal description; we also describe how to regularize based on OM KL divergence.
@@
Figure 5: The true reward achieved by policies regularized with varying amounts of action distribution or occupancy measure regularization using $\chi^{2}$ and KL divergence. The x-axis is the regularization coefficient $\lambda$ normalized by the standard deviation of proxy rewards under the base policy. Dots indicate the median reward and the shaded area is the range over random seeds. For RLHF, AD and OM regularization are equivalent, which is why OM regularization results are not shown for that column.
Experimental setup.  In each of the five environments shown in Figure 2, we train policies with four types of regularization towards the base policy: AD KL, AD $\chi^{2}$, OM KL, and OM $\chi^{2}$. In Appendix A.5, we show that Theorem 5.1 also holds for state-only occupancy measures if the environment’s reward function does not depend on the action; thus, we experiment with regularizing based on both state-action and state-only OM divergence. For each environment and type of regularization, we test a number of regularization coefficients $\lambda$. Theorem 5.1 suggests setting $\lambda=\sigma_{\tilde{R}}\sqrt{1-r^{2}}$ for an $r$-correlated proxy, so for $\chi^{2}$ regularization we test a range of values $\lambda=c\,\sigma_{\tilde{R}}$ from $c=1$ to $10^{-2}$ ($10^{-4}$ for RLHF). Since it is less clear theoretically how to set the coefficient $\lambda$ for KL regularization, we experiment with a wider range of values.
@@
Results.  The results of our experiments are shown in Table 1 and Figure 5. Table 1 shows the median true reward with the best coefficient for each type of regularization. We find that OM regularization consistently outperforms AD regularization across the four non-RLHF environments. In two environments (glucose monitoring and pandemic mitigation), AD regularization fails to improve on the base policy’s true reward at all. Furthermore, $\chi^{2}$ regularization tends to perform similarly to or better than KL regularization across all environments. In RLHF in particular, $\chi^{2}$ regularization leads to a larger improvement over the base policy compared to the industry-standard KL penalty, and is more stable across seeds as well.
@@ Appendix A Proofs and Additional Theoretical Results
$Z(s,a)$ $=\frac{R(s,a)-J({\pi_{\text{base}}},R)}{\sigma_{R}}\qquad\text{and}\qquad\tilde{Z}(s,a)=\frac{\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}$.
@@
$\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$ $\leq\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$
@@
$\quad=\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]+\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$
$\quad=\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right],$
since by definition $\mathbb{E}_{\mu_{\pi_{\text{base}}}}[\tilde{Z}(s,a)]=\mathbb{E}_{\mu_{\pi_{\text{base}}}}[Z(s,a)]=0$. Applying the Cauchy-Schwartz inequality to this difference gives
$\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$
$\quad=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]\Big[\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a)\Big]$
$\quad=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\sqrt{\mu_{\pi_{\text{base}}}(s,a)}\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)\right]\left[\frac{\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a)}{\sqrt{\mu_{\pi_{\text{base}}}(s,a)}}\right]$
$\quad\leq\sqrt{\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{\pi_{\text{base}}}(s,a)\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]^{2}\right)\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{(\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a))^{2}}{\mu_{\pi_{\text{base}}}(s,a)}\right)}$
$\quad=\sqrt{\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)^{2}\right]\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$ (9)
@@
$\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)^{2}\right]$
$\quad=\text{Var}_{\mu_{\pi_{\text{base}}}}\left(\tilde{Z}(s,a)-r\,Z(s,a)\right)$
$\quad=\text{Var}_{\mu_{\pi_{\text{base}}}}\left(\tilde{Z}(s,a)\right)+r^{2}\,\text{Var}_{\mu_{\pi_{\text{base}}}}\left(Z(s,a)\right)-2r\,\text{Cov}_{\mu_{\pi_{\text{base}}}}\left(\tilde{Z}(s,a),Z(s,a)\right)$
@@
using the fact that both reward functions have unit variance under the base policy and that their correlation is $r$. Plugging this into (9) gives the desired result. ∎
A.1.1 Near-Optimal Base Policies
While Theorem 5.1 shows that optimizing the proxy reward with regularization can improve on the base policy’s reward, it does not guarantee that the learned policy will be near-optimal. However, a simple corollary shows that if the base policy is near-optimal, then the learned policy will also be near-optimal.

### v4

@@ 5 Mitigating Reward Hacking with Occupancy Measure Regularization
We now discuss how our new definition of reward hacking motivates better methods for preventing it. Ideally, we would like to be able to optimize a proxy reward function and have it translate into an improvement in true reward over the reference policy. To understand how this might be possible, consider again the traffic environment. The reward hacking policy exhibits behavior which is very unlikely under the reference policy: human drivers hardly ever stop indefinitely on on-ramps. That is, optimizing the proxy reward leads to out-of-distribution states and actions where the correlation that made the proxy good in the first place breaks down. Visually, this can be seen in the second row of Figure 2: the reward hacking policy usually finds state-action pairs with high proxy reward and low true reward that were very unlikely to be reached by the reference policy.
Thus, one solution to prevent reward hacking could be to optimize the proxy reward while avoiding states that are unlikely under the reference policy. The following theorem formalizes this idea.
@@
Suppose that $\tilde{R}$ is an $r$-correlated-proxy for the true reward function $R$, and let $\sigma_{\tilde{R}}$ and $\sigma_{R}$ be defined as in Definition 4.1. Then for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{ref}}}}$ (i.e., $\mu_{{\pi_{\text{ref}}}}(s,a)=0\Rightarrow\mu_{\pi}(s,a)=0$), we have
$\frac{J(\pi,R)-J({\pi_{\text{ref}}},R)}{\sigma_{R}}\geq\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{ref}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}\right),$ (3)
where $\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)=\mathbb{E}_{\mu_{\pi}}\left[\frac{\mu_{\pi}(s,a)}{\mu_{\pi_{\text{ref}}}(s,a)}-1\right]$ is the $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{ref}}}$.
See Appendix A for the proof. Equation (10) gives a lower bound on how much the policy $\pi$ improves over the reference policy $\pi_{\text{ref}}$} in terms of the true reward, normalized by the standard deviation of the true reward under the reference policy. Although this is exactly what we would like to maximize, we must instead optimize the right-hand side (RHS) of (10) since we cannot access the true reward directly. The RHS consists of two terms: one measuring the normalized improvement of the policy $\pi$’s proxy reward over the reference policy and the other penalizing the divergence of $\pi$’s occupancy measure from that of $\pi_{\text{ref}}$}. Optimizing the first term alone often leads to reward hacking, but by both incentivizing $\pi$ to achieve high proxy reward and stay close to the reference policy, $\pi$ can achieve high true reward.
@@
$\text{maximize}\quad J(\pi,\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}\qquad\text{where}\qquad\lambda=\sigma_{\tilde{R}}\sqrt{1-r^{2}}.$ (4)
The amount of regularization needed to improve on the reference policy depends on the strength of correlation $r$. The higher the correlation, the lower the regularization strength $\sqrt{1-r^{2}}$. Given the prefactor of $\frac{1}{r}$ on the RHS of (10), it may appear that a lower correlation leads to a larger gain in true reward. However, Lemma A.2 shows that in fact, the lower bound decreases as a function of $r$.
@@
$\text{maximize}\quad J(\pi,\tilde{R})-\lambda(1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{\text{KL}}\big(\pi(\cdot\mid s_{t})\;\|\;\mu_{\pi_{\text{ref}}}(\cdot\mid s_{t})\big)\right].$ (5)
That is, in RLHF, the expected KL divergence between the action distributions of $\pi$ and $\pi_{\text{ref}}$} is penalized. Action distribution divergence is easy to calculate and optimize, and training LLMs with (5) seems to work well in practice. However, unlike our objective in (4), (5) lacks theoretical guarantees, and it is unclear if it works in other environments.
@@
Figure 4: Our theory also suggests reward hacking can more effectively be prevented by regularizing with $\chi^{2}$ divergence instead of KL divergence. This plot illustrates how $\chi^{2}$ regularization is more effective at preventing reward hacking in RLHF. Both divergences can be written as the expectation of a function $g\big(\log\big(\mu_{\pi}(s,a)/\mu_{{\pi_{\text{ref}}}}(s,a)\big)\big)$ which increases the penalty on state-action pairs based on how far the log-ratio is from zero. The $g(\cdot)$ associated with KL divergence only increases slowly for large log-ratios, so policies trained with KL divergence may produce nonsensical text. In contrast, the $g(\cdot)$ for $\chi^{2}$ divergence increases exponentially, better constraining the LLM to produce text similar to the SFT policy.
Occupancy measure vs. action distribution regularization.  While AD regularization works well for RLHF, this may be because RLHF is essentially a contextual bandit problem, meaning that $\gamma=0$; in this case, OM and AD divergence are equivalent (see Appendix A.3). However, in other cases, AD divergence may not suffice to prevent reward hacking behavior. This is because, in longer-horizon environments, a small change in action distribution at a single state can lead to a much higher probability of reaching undesirable states. Figure 3 shows an example of this in the glucose monitoring environment: policies that are close in action distribution regularization produce vastly different patient glucose levels. Occupancy measure regularization avoids this issue by directly preventing the distribution of glucose levels from differing too much from the reference policy. In Theorem A.5 in the appendix, we show that in general it is impossible to lower bound the improvement in true reward using almost any form of AD regularization, in contrast to our results on OM regularization.
@@
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{{\pi_{\text{ref}}}})=\mathbb{E}_{\mu_{\pi}}\left[d(s,a)+e^{-d(s,a)}\right]\qquad\chi^{2}(\mu_{\pi}\;\|\;\mu_{{\pi_{\text{ref}}}})=\mathbb{E}_{\mu_{\pi}}\left[e^{d(s,a)}+e^{-d(s,a)}\right]$
$\text{where}\qquad d(s,a)=\log\big(\mu_{\pi}(s,a)/\mu_{{\pi_{\text{ref}}}}(s,a)\big).$ (6)
As $d(s,a)$ increases, the optimized policy is visiting state-action pairs that are less likely under the reference policy. However, KL divergence only penalizes $d(s,a)$ linearly, while $\chi^{2}$ penalizes it exponentially, resulting in stronger regularization even with a low coefficient. Figure 4 plots the functions in (6) and shows how in practice $\chi^{2}$ divergence better prevents reward hacking in RLHF.
@@
Table 1: We compare using various types of regularization to prevent reward hacking in the five environments form Figure 2. The median true reward and standard deviation across 5 random seeds is shown for the best regularization coefficient for each type of regularization. The bottom rows show results for the baselines: the reference policy $\pi_{\text{ref}}$}, a policy trained on the proxy reward without regularization (exhibiting reward hacking), and a policy trained on the true reward function (impossible in practice, but included as an upper bound on performance). We find that occupancy measure regularization consistently improves on action distribution regularization, and that $\chi^{2}$ divergence often outperforms KL divergence.
@@ 6 Experiments
$\text{and}\qquad\phi$ $=\arg\min_{\phi^{\prime}}\;\mathbb{E}_{\mu_{\pi}}\Big[\,\log(1+e^{-\hat{d}_{\phi^{\prime}}(s,a)})\,\Big]+\,\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\Big[\,\log(1+e^{\hat{d}_{\phi^{\prime}}(s,a)})\,\Big].$ (7)
That is, optimizing a discriminator network $\hat{d}_{\phi}$ to minimize the given loss can be used to estimate and optimize $\chi^{2}$ OM divergence. We alternately train $\hat{d}_{\phi}$ via gradient descent on (7) and the policy $\pi$ via PPO based on the adjusted reward. We call this algorithm Occupancy-Regularized Policy Optimization (ORPO). See Appendix B for a full derivation of these approximations and Algorithm 1 for a formal description; we also describe how to regularize based on OM KL divergence.
@@
Figure 5: The true reward achieved by policies regularized with varying amounts of action distribution or occupancy measure regularization using $\chi^{2}$ and KL divergence. The x-axis is the regularization coefficient $\lambda$ normalized by the standard deviation of proxy rewards under the reference policy. Dots indicate the median reward and the shaded area is the range over random seeds. For RLHF, AD and OM regularization are equivalent, which is why OM regularization results are not shown for that column.
Experimental setup.  In each of the five environments shown in Figure 2, we train policies with four types of regularization towards the reference policy: AD KL, AD $\chi^{2}$, OM KL, and OM $\chi^{2}$. In Appendix A.5, we show that Theorem 5.1 also holds for state-only occupancy measures if the environment’s reward function does not depend on the action; thus, we experiment with regularizing based on both state-action and state-only OM divergence. For each environment and type of regularization, we test a number of regularization coefficients $\lambda$. Theorem 5.1 suggests setting $\lambda=\sigma_{\tilde{R}}\sqrt{1-r^{2}}$ for an $r$-correlated proxy, so for $\chi^{2}$ regularization we test a range of values $\lambda=c\,\sigma_{\tilde{R}}$ from $c=1$ to $10^{-2}$ ($10^{-4}$ for RLHF). Since it is less clear theoretically how to set the coefficient $\lambda$ for KL regularization, we experiment with a wider range of values.
@@
Results.  The results of our experiments are shown in Table 1 and Figure 5. Table 1 shows the median true reward with the best coefficient for each type of regularization. We find that OM regularization consistently outperforms AD regularization across the four non-RLHF environments. In two environments (glucose monitoring and pandemic mitigation), AD regularization fails to improve on the reference policy’s true reward at all. Furthermore, $\chi^{2}$ regularization tends to perform similarly to or better than KL regularization across all environments. In RLHF in particular, $\chi^{2}$ regularization leads to a larger improvement over the reference policy compared to the industry-standard KL penalty, and is more stable across seeds as well.
@@ Appendix A Proofs and Additional Theoretical Results
$Z(s,a)$ $=\frac{R(s,a)-J({\pi_{\text{ref}}},R)}{\sigma_{R}}\qquad\text{and}\qquad\tilde{Z}(s,a)=\frac{\tilde{R}(s,a)-J({\pi_{\text{ref}}},\tilde{R})}{\sigma_{\tilde{R}}}$.
Combining (1) and (2), we can see that the return of a policy can be re-written as the expected reward for states and actions sampled from the occupancy measure:
@@
$\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$ $\leq\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}.$
@@
$\quad=\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]+\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$
$\quad=\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right],$
since by definition $\mathbb{E}_{\mu_{\pi_{\text{ref}}}}[\tilde{Z}(s,a)]=\mathbb{E}_{\mu_{\pi_{\text{ref}}}}[Z(s,a)]=0$. Applying the Cauchy-Schwartz inequality to this difference gives
$\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$
$\quad=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]\Big[\mu_{\pi}(s,a)-\mu_{\pi_{\text{ref}}}(s,a)\Big]$
$\quad=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\sqrt{\mu_{\pi_{\text{ref}}}(s,a)}\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)\right]\left[\frac{\mu_{\pi}(s,a)-\mu_{\pi_{\text{ref}}}(s,a)}{\sqrt{\mu_{\pi_{\text{ref}}}(s,a)}}\right]$
$\quad\leq\sqrt{\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{\pi_{\text{ref}}}(s,a)\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]^{2}\right)\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{(\mu_{\pi}(s,a)-\mu_{\pi_{\text{ref}}}(s,a))^{2}}{\mu_{\pi_{\text{ref}}}(s,a)}\right)}$
$\quad=\sqrt{\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)^{2}\right]\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{ref}}}\right)}.$ (9)
@@
$\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)^{2}\right]$
$\quad=\text{Var}_{\mu_{\pi_{\text{ref}}}}\left(\tilde{Z}(s,a)-r\,Z(s,a)\right)$
$\quad=\text{Var}_{\mu_{\pi_{\text{ref}}}}\left(\tilde{Z}(s,a)\right)+r^{2}\,\text{Var}_{\mu_{\pi_{\text{ref}}}}\left(Z(s,a)\right)-2r\,\text{Cov}_{\mu_{\pi_{\text{ref}}}}\left(\tilde{Z}(s,a),Z(s,a)\right)$
@@
using the fact that both reward functions have unit variance under the reference policy and that their correlation is $r$. Plugging this into (9) gives the desired result. ∎
A.1.1 Near-Optimal reference policies
While Theorem 5.1 shows that optimizing the proxy reward with regularization can improve on the reference policy’s reward, it does not guarantee that the learned policy will be near-optimal. However, a simple corollary shows that if the reference policy is near-optimal, then the learned policy will also be near-optimal.